OpenStudy (anonymous):

2^(x + 3) = 7. I ended up with x = 20.2534966 but it wasn't one of the answer choices

OpenStudy (anonymous):

What were the choices? Were they exact or decimal?

OpenStudy (anonymous):

x ≈ -5.19 x ≈ 0.94 x ≈ -0.19 x ≈ 2.81

OpenStudy (anonymous):

\[x=\log_{2} (7)-3=\frac{\ln 7}{\ln 2}-3\]

OpenStudy (anonymous):

oh. How do i get that? i got (x+3)log 2 = 7 lol

OpenStudy (lgbasallote):

it should be \[(x+3)\log 2 = \log 7\] if you're going that way

OpenStudy (lgbasallote):

then divide both sides by \(\log 2\) \[x+ 3 = \frac{\log 7}{\log 2}\] \[x = \frac{log 7}{\log 2} - 3\]

OpenStudy (anonymous):

oh so you're supposed to single out the x if there's only one in the equation?

OpenStudy (lgbasallote):

yup. and also because you're looking for x

OpenStudy (lgbasallote):

wow what a coincidence...the result is one of the choices ;)

OpenStudy (anonymous):

ah i see. alright thanks :D